Write the standard form of the equation of the parabola with the given focus (-1,0) and vertex of (0,0). Then state the directrix.
Also with the vertex (2,3) and given focus (-2,0)

Answer by lwsshak3(11628)

Write the standard form of the equation of the parabola with the given focus (-1,0) and vertex of (0,0). Then state the directrix.
Also with the vertex (2,3) and given focus (-2,0)
***
vertex: (0,0)
focus: (-1,0)
axis of symmetry: y=0
parabola opens leftward:
Its basic equation: (y-k)^2=-4p(x-h), (h,k)=(x,y) coordinates of the vertex
p=1(distance from focus to vertex on the axis of symmetry)
4p=4
directrix: x=1
Equation: y^2=-4x
..
vertex: (2,3)
focus: (-2,3)
axis of symmetry: y=3
parabola opens leftward:
Its basic equation: (y-k)^2=-4p(x-h), (h,k)=(x,y) coordinates of the vertex
p=4(distance from focus to vertex on the axis of symmetry)
4p=16
directrix: x=6
Equation: (y-3)^2=-16(x-2)